# Morita equivalence of Nijenhuis structures

### v1

@@ 3.2 The Lie correspondence for (1,1)-tensor fields
Let $\mathscr{G}$ be a source-simply connected Lie groupoid. The Lie functor establishes a bijection between multiplicative $(1,1)$-tensor fields $K\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ on $\mathscr{G}$ and IM $(1,1)$-tensor fields $K_{*}\in\Omega^{1}(A_{\mathscr{G}},TA_{\mathscr{G}})$ on $A_{\mathscr{G}}$. Moreover, $K$ is Nijenhuis if and only if its correspoding $K_{*}$ is Nijenhuis.
In particular, there is a one-to-one correspondence between (source-simply connected) Nijenhuis groupoids and (integrable) infinitesimal Nijenhuis structures. The IM $1$-derivation $\mathscr{D}=\left(\nabla,\ell,r\right)$ on $A_{\mathscr{G}}$ corresponding to a multiplicative $(1,1)$-tensor field $K\in\Omega^{1}\left(\mathscr{G},T\mathscr{G}\right)$ is defined by the equations
$K\left(\overrightarrow{\xi}\right)=\overrightarrow{\ell(\xi)},\quad\mathrm{d}\mathbf{t}\left(K(v)\right)=r\left(\mathrm{d}\mathbf{t}(v)\right)\quad\text{and}\quad\left(\mathscr{L}_{\overrightarrow{\xi}}K\right)(v)=\overrightarrow{\nabla_{\mathrm{d}\mathbf{t}(v)}(\xi)},$ (24)
@@
Integration of tangent lifts.
The tangent bundle $TP\to P$ of a manifold $P$, equipped with its canonical Lie algebroid structure, integrates to the pair groupoid $\mathrm{Pair}(P)=(P\times P\rightrightarrows P)$. The tangent lift $J^{\mathrm{tg}}\in\Omega^{1}(TP,T(TP))$ of a vector-valued $1$-form $J\in\Omega^{1}(P,TP)$ is a $(1,1)$-tensor field on $TP$ given by the conjugation of $\mathrm{d}J$ by the canonical involution $\kappa_{P}:T(TP)\to T(TP)$; that is,
$J^{\mathrm{tg}}\coloneqq\kappa_{P}\circ\mathrm{d}J\circ\kappa_{P}:T(TP)\to T(TP).$ (25)
The tangent lift $J^{\mathrm{tg}}$ is an IM $(1,1)$-tensor field on $TP$. We now focus on its integration to a multiplicative $(1,1)$-tensor field on $\mathrm{Pair}(P)$, because it is used in the integration of Morita equivalences.
@@
We next show that $J^{\mathrm{tg}}$ is the differentiation of $J\times J$ under the Lie functor. Theorem 7.1 of [26] establishes a canonical isomorphism between the double vector bundles $T\left(A_{\mathscr{G}}\right)$ and $A_{T\mathscr{G}}$ for any Lie groupoid $\mathscr{G}$, which is an isomorphism of Lie algebroids. This isomorphism is the restriction of the canonical involution $\kappa_{G_{1}}:T(TG_{1})\to T(TG_{1})$ to a map between $TA_{\mathscr{G}}=T\left(\mathbf{u}_{\mathscr{G}}^{*}\operatorname{ker}\mathrm{d}\mathbf{s}_{\mathscr{G}}\right)$ and $A_{T\mathscr{G}}=\mathbf{u}_{T\mathscr{G}}^{*}\operatorname{ker}\mathrm{d}\mathbf{s}_{T\mathscr{G}}$. In the present case, $G_{1}=P\times P$ and $G_{0}=P$, whereas $A_{\mathscr{G}}=TP\to P$ is the tangent bundle of $P$ and $A_{T\mathscr{G}}=T(TP)\to TP$ is the tangent Lie algebroid. It follows that $\kappa_{G_{1}}=\kappa_{P}\times\kappa_{P}$, and the isomorphism corresponds to the upper arrow in the following diagram:
@@
Every IM $(1,1)$-tensor field $R\in\Omega^{1}(TP,T(TP))$ on a tangent bundle $TP\to P$ is of the form $R=J^{\mathrm{tg}}$ for some $J\in\Omega^{1}(P,TP)$ (see [16, Corollary 3.5]).
3.3 Morita equivalence of infinitesimal structures
In this subsection, we introduce a notion of Morita equivalence for IM $(1,1)$-tensor fields on Lie algebroids. While distinct notions of Morita equivalence for Lie algebroids exist in the literature—defined, for instance, via isomorphic pullback Lie algebroids [20] or in terms of preserved geometric data [37]—we adopt a simplified definition (Definition 3.8) resulting from differentiating the global formulation of Morita equivalence via principal bibundles. This approach is analogous to the construction of Morita equivalence for Poisson manifolds, which is obtained by differentiating Morita equivalences of symplectic groupoids [39]. Consequently, our framework is built upon infinitesimal actions. Because this differentiation yields complete actions, this definition applies exclusively to integrable Lie algebroids (see Remark 3.9), as is the case for Poisson structures.
@@ 3.3 Morita equivalence of infinitesimal structures
$\rho(\xi)=\mathrm{d}\mu\left(\mathscr{a}(\xi)\right)\qquad\text{and}\qquad\mathscr{a}\left(f\xi\right)=(\mu^{*}f)\mathscr{a}(\xi),$ (26)
@@
    $\left[\mathscr{a}_{1}(\xi),\mathscr{a}_{2}(\zeta)\right]=0.$ (27)
@@
    $\operatorname{im}\left(\mathscr{a}_{1\,p}\right)=\operatorname{ker}\mathrm{d}\mu_{2\,p}\qquad\text{and}\qquad\operatorname{im}\left(\mathscr{a}_{2\,p}\right)=\operatorname{ker}\mathrm{d}\mu_{1\,p}.$ (28)
@@
The existence of a complete action of a Lie algebroid on a surjective submersion is equivalent to the integrability of the algebroid ([1, Theorem 4.1] and [11, Theorem 8]). Consequently, Definition 3.8 applies exclusively to integrable Lie algebroids. This ensures that infinitesimal Morita equivalences of Lie algebroids correspond to Morita equivalences of the respective source-simply connected Lie groupoids (Lemma 3.10). In contrast, the notion of Morita equivalence in [20] is defined by requiring the pullback Lie algebroids $\mu_{1}^{!}A_{1}$ and $\mu_{2}^{!}A_{2}$ to be isomorphic (with simply-connected $\mu_{i}$-fibers), a condition that arises from differentiating Morita morphisms rather than bibundles.
It is known that a complete action of a Lie algebroid integrates to a global action of the corresponding source-simply connected Lie groupoid (see [33, Theorem 2.1] and [31, Theorem 5.3]). More precisely, if $\mathscr{a}:\Gamma(A)\to\mathfrak{X}(P)$ is a complete action of $A\to M$ on $\mu:P\to M$, then there exists a unique action $\mathscr{A}:G\times_{(\mathbf{s},\mu)}P\to P$ of the source-simply connected Lie groupoid $G\rightrightarrows M$ integrating $A$. The infinitesimal and global actions are related by
$\mathscr{a}(\xi)_{p}=\left.\frac{\mathrm{d}}{\mathrm{d}\tau}\right|_{\tau=0}\mathscr{A}(\varphi(\tau),p),$ (29)
@@
The sum of the actions.
The conditions on the Morita equivalence for Lie algebroids allow us to construct an action of the product algebroid $A_{1}\times A_{2}^{\mathrm{op}}$ on the submersion $\mu\coloneqq(\mu_{1},\mu_{2}):P\to M_{1}\times M_{2}$. Consider the map
$\Gamma(A_{1})\oplus\Gamma(A_{2}^{\mathrm{op}})\to\mathfrak{X}(P),\qquad(\xi_{1},\xi_{2})\mapsto\mathscr{a}_{1}(\xi_{1})+\mathscr{a}_{2}(\xi_{2}).$ (30)
Extending this map by $C^{\infty}(M_{1}\times M_{2})$-linearity yields a Lie algebroid morphism $\mathscr{a}:\Gamma(A_{1}\times A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$. We show next that $\mathscr{a}$ defines an action on $\mu:P\to M_{1}\times M_{2}$. Since the actions commute,
$\mathscr{a}\left(\left[(\xi_{1},\zeta_{1}),(\xi_{2},\zeta_{2})\right]_{A_{1}\times A_{2}^{\mathrm{op}}}\right)$ $=\mathscr{a}_{1}\left(\left[\xi_{1},\xi_{2}\right]_{A_{1}}\right)+\mathscr{a}_{2}\left(\left[\zeta_{1},\zeta_{2}\right]_{A_{2}}\right)$
$=\left[\mathscr{a}_{1}(\xi_{1})+\mathscr{a}_{2}(\zeta_{1}),\mathscr{a}_{1}(\xi_{2})+\mathscr{a}_{2}(\zeta_{2})\right]$
$=\left[\mathscr{a}\left(\xi_{1},\zeta_{1}\right),\mathscr{a}\left(\xi_{2},\zeta_{2}\right)\right].$
Furthermore, using conditions (28), we obtain
$\mathrm{d}\mu\left(\mathscr{a}\left(\xi,\zeta\right)\right)$ $=\left(\mathrm{d}\mu_{1}\left(\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)\right),\mathrm{d}\mu_{2}\left(\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)\right)\right)$
$=\left(\mathrm{d}\mu_{1}\left(\mathscr{a}_{1}(\xi)\right),\mathrm{d}\mu_{2}\left(\mathscr{a}_{2}(\zeta)\right)\right)$
$=\rho_{A_{1}\times A_{2}^{\mathrm{op}}}\left(\xi,\zeta\right).$
We call $\mathscr{a}:\Gamma(A_{1}\times A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$ the sum of the actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$.
Pullback of linear $(1,1)$-tensors.
We introduce the pullback of linear $(1,1)$-tensor fields, which is required to define Morita equivalence of IM $(1,1)$-tensor fields. Let $E\to M$ be a vector bundle and $R:TE\to TE$ a linear $(1,1)$-tensor field on $E$ covering a $(1,1)$-tensor field $r:TM\to TM$. Let $\mu:P\to M$ be a smooth map and let $J:TP\to TP$ be a $(1,1)$-tensor field on $P$ which is $\mu$-related to $r$.
@@
$\mathrm{d}\mathscr{a}\circ\mu^{*}(R_{1}\times R_{2})=J^{\mathrm{tg}}\circ\mathrm{d}\mathscr{a},$ (32)
where $\mathscr{a}=\mathscr{a}_{1}+\mathscr{a}_{2}$ denotes the sum of the actions and $\mu\coloneqq(\mu_{1},\mu_{2})$. This relation is called a Morita equivalence of Nijenhuis IM tensor fields if $R_{1}$, $R_{2}$, and $J$ are Nijenhuis tensor fields.
@@
$R_{1}\times R_{2}\times J^{\mathrm{tg}}\left(T\mathrm{Graph}(\mathscr{a})\right)\subseteq T\mathrm{Graph}(\mathscr{a}).$ (33)
@@
Consider two holomorphic Lie algebroids, $(A_{1},I_{1})$ and $(A_{2},I_{2})$, and suppose they are Morita equivalent via a bibundle $M_{1}\xleftarrow{\ \mu_{1}\ }(P,J)\xrightarrow{\ \mu_{2}\ }M_{2}$, where $(P,J)$ is a complex manifold. One can verify that the tangent lift of $J$ is a complex structure on $TP$. Consequently, equation (32) guarantees that the infinitesimal action $\mathscr{a}:\Gamma(A_{1}\times A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$ is holomorphic.
Infinitesimal Morita equivalence in terms of $1$-derivations.
We now translate the formulation of Morita equivalence of IM tensor fields into the algebraic language of $1$-derivations. An advantage of this formulation is that it avoids the need to explicitly perform the differential of the infinitesimal actions by separating the tensorial conditions from the differential ones. As seen below, Equations (34) and (35) are purely tensorial; the differential content is entirely isolated within condition (36) for the connection-like components.
To fix notation for the following definition, let $\mathscr{D}_{1}=(\nabla^{1},\ell_{1},r_{1})$ and $\mathscr{D}_{2}=(\nabla^{2},\ell_{2},r_{2})$ be IM $1$-derivations on the Lie algebroids $A_{1}\to M_{1}$ and $A_{2}\to M_{2}$, respectively. Additionally, let $\mathscr{a}_{1}:\Gamma(A_{1})\to\mathfrak{X}(P)$ and $\mathscr{a}_{2}:\Gamma(A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$ denote the infinitesimal actions, and set $\mathscr{a}\coloneqq\mathscr{a}_{1}+\mathscr{a}_{2}$.
@@
A Morita equivalence of $1$-derivations $\mathscr{D}_{1}=(\nabla^{1},\ell_{1},r_{1})$ and $\mathscr{D}_{2}=(\nabla^{2},\ell_{2},r_{2})$ consists of a Morita $(A_{1},A_{2})$-bibundle $M_{1}\xleftarrow{\ \mu_{1}\ }P\xrightarrow{\ \mu_{2}\ }M_{2}$, together with a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$, satisfying
$\mathrm{d}\mu\circ J=(r_{1}\times r_{2})\circ\mathrm{d}\mu$ (34)
$J\circ\mathscr{a}=\mathscr{a}\circ(\ell_{1}\times\ell_{2})$ (35)
$\nabla_{v}^{J}\left(\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)\right)=\mathscr{a}_{1}\left(\nabla^{1}_{\mathrm{d}\mu_{1}(v)}(\xi)\right)+\mathscr{a}_{2}\left(\nabla^{2}_{\mathrm{d}\mu_{2}(v)}(\zeta)\right)$ (36)
@@
A Morita equivalence of Lie algebroids $M_{1}\xleftarrow{\ \mu_{1}\ }(P,J)\xrightarrow{\ \mu_{2}\ }M_{2}$ equipped with a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$ is a Morita equivalence of IM $(1,1)$-tensor fields $R_{1}\in\Omega^{1}(A_{1},TA_{1})$ and $R_{2}\in\Omega^{1}(A_{2},TA_{2})$ if and only if it is a Morita equivalence of the respective $1$-derivations $\mathscr{D}_{1}$ and $\mathscr{D}_{2}$.
The proof of this proposition relies on the following lemma, which characterizes the pullback of $1$-derivations.
@@
Let $\pi:E\to M$ be a vector bundle equipped with a linear $(1,1)$-tensor field $R\in\Omega^{1}(E,TE)$ covering $r\in\Omega^{1}(M,TM)$, and let $\mathscr{D}=(\nabla,\ell,r)$ be its associated $1$-derivation. Let $\mu:P\to M$ be a smooth map and suppose $J\in\Omega^{1}(P,TP)$ is a $(1,1)$-tensor field $\mu$-related to $r$. Then the $1$-derivation associated with the pullback $\mu^{*}R$ is given by the triple $\mu^{*}\mathscr{D}=\left(\mu^{*}\nabla,\mu^{*}\ell,J\right)$, whose components are uniquely determined by their action on the pullback of sections:
$(\mu^{*}\nabla)_{v}(\mu^{*}\xi)=\mu^{*}\left(\nabla_{\mathrm{d}\mu(v)}\xi\right),$ (37)
$(\mu^{*}\ell)(\mu^{*}\xi)=\mu^{*}(\ell(\xi)),$ (38)
@@
Now, applying the correspondence (20) between linear $(1,1)$-tensors and $1$-derivations, we directly identify the base component as $J\in\Omega^{1}(P,TP)$. To check the other components, note that the vertical lift of a pullback section $\mu^{*}\xi\in\Gamma(\mu^{*}E)$ at $(p,e)\in\mu^{*}E$ is given by:
@@
On the other hand, since the projection of $(v,w)$ to $TP$ is $\mathrm{d}(\mu^{*}\pi)(v,w)=v$, the definition of the connection via the correspondence (20) yields $\left((\mu^{*}\nabla)_{v}(\mu^{*}\xi)\right)^{\uparrow}=\mathscr{L}_{(\mu^{*}\xi)^{\uparrow}}(\mu^{*}R)(v,w)$. Thus, we conclude:
@@
We now proceed to the proof of Proposition 3.16. To this end, we recall the characterization of morphisms of related linear tensor fields established in [8, Theorem 2.1]. Let $R_{i}\in\Omega^{1}(E_{i},TE_{i})$ be linear $(1,1)$-tensor fields on vector bundles $E_{i}\to M$ for $i=1,2$, and let $\Phi:E_{1}\to E_{2}$ be a vector bundle map covering the identity. Then
$\mathrm{d}\Phi\circ R_{1}=R_{2}\circ\mathrm{d}\Phi\qquad\iff\qquad\Phi\circ\nabla_{v}^{1}=\nabla_{v}^{2}\circ\Phi,\quad\Phi\circ\ell_{1}=\ell_{2}\circ\Phi\quad\text{and}\quad r_{1}=r_{2},$ (39)

### v2

@@ 3.2. The Lie correspondence for (1,1)-tensor fields
Let $\mathscr{G}$ be a source-simply connected Lie groupoid. The Lie functor establishes a bijection between multiplicative $(1,1)$-tensor fields $K\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ on $\mathscr{G}$ and IM $(1,1)$-tensor fields $R_{K}\in\Omega^{1}(A_{\mathscr{G}},TA_{\mathscr{G}})$ on $A_{\mathscr{G}}$. Moreover, $K$ is Nijenhuis if and only if its corresponding $R_{K}$ is Nijenhuis.
In particular, there is a one-to-one correspondence between (source-simply connected) Nijenhuis groupoids and (integrable) infinitesimal Nijenhuis structures. The IM 11-derivation $\mathscr{D}=\left(\nabla,\ell,r\right)$ on $A_{\mathscr{G}}$ corresponding to a multiplicative $(1,1)$-tensor field $K\in\Omega^{1}\left(\mathscr{G},T\mathscr{G}\right)$ is defined by the equations
(24) $K\left(\overrightarrow{\xi}\right)=\overrightarrow{\ell(\xi)},\quad\mathrm{d}\mathbf{t}\left(K(v)\right)=r\left(\mathrm{d}\mathbf{t}(v)\right)\quad\text{and}\quad\left(\mathscr{L}_{\overrightarrow{\xi}}K\right)(v)=\overrightarrow{\nabla_{\mathrm{d}\mathbf{t}(v)}(\xi)},$
@@
3.2.1. Integration of tangent lifts
The tangent bundle $TP\to P$ of a manifold $P$, equipped with its canonical Lie algebroid structure, integrates to the pair groupoid $\mathrm{Pair}(P)=(P\times P\rightrightarrows P)$. The tangent lift $J^{\mathrm{tg}}\in\Omega^{1}(TP,T(TP))$ of a vector-valued 11-form $J\in\Omega^{1}(P,TP)$ is a $(1,1)$-tensor field on TPTP given by the conjugation of $\mathrm{d}J$ by the canonical involution $\kappa_{P}:T(TP)\to T(TP)$; that is,
(25) $J^{\mathrm{tg}}\coloneqq\kappa_{P}\circ\mathrm{d}J\circ\kappa_{P}:T(TP)\to T(TP).$
The tangent lift $J^{\mathrm{tg}}$ is an IM $(1,1)$-tensor field on TPTP. We now focus on its integration to a multiplicative $(1,1)$-tensor field on $\mathrm{Pair}(P)$, because it is used in the integration of Morita equivalences.
@@ 3.2.1. Integration of tangent lifts
We next show that $J^{\mathrm{tg}}$ is the differentiation of $J\times J$ under the Lie functor. Theorem 7.1 of [undefaa] establishes a canonical isomorphism between the double vector bundles $T\left(A_{\mathscr{G}}\right)$ and $A_{T\mathscr{G}}$ for any Lie groupoid $\mathscr{G}$, which is an isomorphism of Lie algebroids. This isomorphism is the restriction of the canonical involution $\kappa_{G_{1}}:T(TG_{1})\to T(TG_{1})$ to a map between $TA_{\mathscr{G}}=T\left(\mathbf{u}_{\mathscr{G}}^{*}\operatorname{ker}\mathrm{d}\mathbf{s}_{\mathscr{G}}\right)$ and $A_{T\mathscr{G}}=\mathbf{u}_{T\mathscr{G}}^{*}\operatorname{ker}\mathrm{d}\mathbf{s}_{T\mathscr{G}}$. In the present case, $G_{1}=P\times P$ and $G_{0}=P$, whereas $A_{\mathscr{G}}=TP\to P$ is the tangent bundle of $P$ and $A_{T\mathscr{G}}=T(TP)\to TP$ is the tangent Lie algebroid. It follows that $\kappa_{G_{1}}=\kappa_{P}\times\kappa_{P}$, and the isomorphism corresponds to the upper arrow in the following diagram:
@@
Every IM $(1,1)$-tensor field $R\in\Omega^{1}(TP,T(TP))$ on a tangent bundle $TP\to P$ is of the form $R=J^{\mathrm{tg}}$ for some $J\in\Omega^{1}(P,TP)$ (see [undefp, Corollary 3.5]).
$3.$ Morita equivalence of infinitesimal structures
In this subsection, we introduce a notion of Morita equivalence for IM $(1,1)$-tensor fields on Lie algebroids. While distinct notions of Morita equivalence for Lie algebroids exist in the literature—defined, for instance, via isomorphic pullback Lie algebroids [undefs] or in terms of preserved geometric data [undefak]—we adopt a simplified definition (Definition 3.8) resulting from differentiating the global formulation of Morita equivalence via principal bibundles. This approach is analogous to the construction of Morita equivalence for Poisson manifolds, which is obtained by differentiating Morita equivalences of symplectic groupoids [undefan]. Consequently, our framework is built upon infinitesimal actions. Because this differentiation yields complete actions, this definition applies exclusively to integrable Lie algebroids (see Remark 3.9), as is the case for Poisson structures.
@@
(26) $\rho(\xi)=\mathrm{d}\mu\left(\mathscr{a}(\xi)\right)\qquad\text{and}\qquad\mathscr{a}\left(f\xi\right)=(\mu^{*}f)\mathscr{a}(\xi),$
@@
    (27) $\left[\mathscr{a}_{1}(\xi),\mathscr{a}_{2}(\zeta)\right]=0.$
@@
    (28) $\operatorname{im}\left(\mathscr{a}_{1\,p}\right)=\operatorname{ker}\mathrm{d}\mu_{2\,p}\qquad\text{and}\qquad\operatorname{im}\left(\mathscr{a}_{2\,p}\right)=\operatorname{ker}\mathrm{d}\mu_{1\,p}.$
@@
The existence of a complete action of a Lie algebroid on a surjective submersion is equivalent to the integrability of the algebroid ([undefa, Theorem 4.1] and [undefl, Theorem 8]). Consequently, Definition 3.8 applies exclusively to integrable Lie algebroids. This ensures that infinitesimal Morita equivalences of Lie algebroids correspond to Morita equivalences of the respective source-simply connected Lie groupoids (Lemma 3.11). In contrast, the notion of Morita equivalence in [undefs] is defined by requiring the pullback Lie algebroids $\mu_{1}^{!}A_{1}$ and $\mu_{2}^{!}A_{2}$ to be isomorphic (with simply-connected $\mu_{i}$-fibers), a condition that arises from differentiating Morita morphisms rather than bibundles.
Remark 3.10.
The conditions for Morita equivalence of Lie algebroids allow us to construct an action of the product algebroid $A_{1}\times A_{2}^{\mathrm{op}}$ on the submersion $\mu\coloneqq(\mu_{1},\mu_{2})\colon P\to M_{1}\times M_{2}$. This action is defined by
(29) $\Gamma(A_{1})\oplus\Gamma(A_{2}^{\mathrm{op}})\to\mathfrak{X}(P),\qquad(\xi_{1},\xi_{2})\mapsto\mathscr{a}_{1}(\xi_{1})+\mathscr{a}_{2}(\xi_{2}),$
and extended by $C^{\infty}(M_{1}\times M_{2})$-linearity to a Lie algebroid morphism $\mathscr{a}\colon\Gamma(A_{1}\times A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$. We refer to this morphism as the total (infinitesimal) action, and we denote it by $\mathscr{a}=\mathscr{a}_{1}+\mathscr{a}_{2}$.
It is known that a complete action of a Lie algebroid integrates to a global action of the corresponding source-simply connected Lie groupoid (see [undefag, Theorem 2.1] and [undefaf, Theorem 5.3]). More precisely, if $\mathscr{a}:\Gamma(A)\to\mathfrak{X}(P)$ is a complete action of $A\to M$ on $\mu:P\to M$, then there exists a unique action $\mathscr{A}:G\times_{(\mathbf{s},\mu)}P\to P$ of the source-simply connected Lie groupoid $G\rightrightarrows M$ integrating AA. The infinitesimal and global actions are related by
(30) $\mathscr{a}(\xi)_{p}=\left.\frac{\mathrm{d}}{\mathrm{d}\tau}\right|_{\tau=0}\mathscr{A}(\varphi(\tau),p),$
@@
3.3.1. Infinitesimal Morita equivalence of linear $(1,1)$-tensor fields
To define the Morita equivalence of IM $(1,1)$-tensor fields, we must first introduce the pullback of linear $(1,1)$-tensor fields. Let $E\to M$ be a vector bundle, and let $R\colon TE\to TE$ be a linear $(1,1)$-tensor field on $E$ covering a $(1,1)$-tensor field $r\colon TM\to TM$. Furthermore, let $\mu\colon P\to M$ be a smooth map, and let $J\colon TP\to TP$ be a $(1,1)$-tensor field on $P$ that is $\mu$-related to $r$.
@@ 3.3.1. Infinitesimal Morita equivalence of linear $(1,1)$-tensor fields
(32) $\mathrm{d}\mathscr{a}\circ\mu^{*}(R_{1}\times R_{2})=J^{\mathrm{tg}}\circ\mathrm{d}\mathscr{a},$
where $\mathscr{a}=\mathscr{a}_{1}+\mathscr{a}_{2}$ denotes the total action and $\mu\coloneqq(\mu_{1},\mu_{2})$. This relation is called a Morita equivalence of Nijenhuis IM tensor fields if $R_{1}$, $R_{2}$, and $J$ are Nijenhuis tensor fields.
@@
(33) $R_{1}\times R_{2}\times J^{\mathrm{tg}}\left(T\mathrm{Graph}(\mathscr{a})\right)\subseteq T\mathrm{Graph}(\mathscr{a}).$
@@
Consider two holomorphic Lie algebroids, $(A_{1},I_{1})$ and $(A_{2},I_{2})$, and suppose they are Morita equivalent via a bibundle $M_{1}\xleftarrow{\ \mu_{1}\ }(P,J)\xrightarrow{\ \mu_{2}\ }M_{2}$, where $(P,J)$ is a complex manifold. One can verify that the tangent lift of $J$ is a complex structure on TPTP. Consequently, Equation (32) guarantees that the infinitesimal action $\mathscr{a}:\Gamma(A_{1}\times A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$ is holomorphic.
3.3.2. Infinitesimal Morita equivalence of 11-derivations.
We now translate the formulation of Morita equivalence of IM tensor fields into the algebraic language of 11-derivations. An advantage of this formulation is that it avoids the need to explicitly perform the differential of the infinitesimal actions by separating the tensorial conditions from the differential ones. As seen below, Equations (34) and (35) are purely tensorial; the differential content is entirely isolated within condition (36) for the connection-like components.
To fix notation for the following definition, let $\mathscr{D}_{1}=(\nabla^{1},\ell_{1},r_{1})$ and $\mathscr{D}_{2}=(\nabla^{2},\ell_{2},r_{2})$ be IM 11-derivations on the Lie algebroids $A_{1}\to M_{1}$ and $A_{2}\to M_{2}$, respectively. Additionally, let $\mathscr{a}_{1}:\Gamma(A_{1})\to\mathfrak{X}(P)$ and $\mathscr{a}_{2}:\Gamma(A_{2}^{\mathrm{op}})\to\mathfrak{X}(P)$ denote the infinitesimal actions, and set $\mathscr{a}\coloneqq\mathscr{a}_{1}+\mathscr{a}_{2}$.
@@
A Morita equivalence of 11-derivations $\mathscr{D}_{1}=(\nabla^{1},\ell_{1},r_{1})$ and $\mathscr{D}_{2}=(\nabla^{2},\ell_{2},r_{2})$ consists of a Morita $(A_{1},A_{2})$-bibundle $M_{1}\xleftarrow{\ \mu_{1}\ }P\xrightarrow{\ \mu_{2}\ }M_{2}$, together with a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$, satisfying
(34) $\mathrm{d}\mu\circ J=(r_{1}\times r_{2})\circ\mathrm{d}\mu$
(35) $J\circ\mathscr{a}=\mathscr{a}\circ(\ell_{1}\times\ell_{2})$
(36) $\nabla_{v}^{J}\left(\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)\right)=\mathscr{a}_{1}\left(\nabla^{1}_{\mathrm{d}\mu_{1}(v)}(\xi)\right)+\mathscr{a}_{2}\left(\nabla^{2}_{\mathrm{d}\mu_{2}(v)}(\zeta)\right)$
@@
A Morita equivalence of Lie algebroids $M_{1}\xleftarrow{\ \mu_{1}\ }(P,J)\xrightarrow{\ \mu_{2}\ }M_{2}$ equipped with a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$ is a Morita equivalence of IM $(1,1)$-tensor fields $R_{1}\in\Omega^{1}(A_{1},TA_{1})$ and $R_{2}\in\Omega^{1}(A_{2},TA_{2})$ if and only if it is a Morita equivalence of the respective 11-derivations $\mathscr{D}_{1}$ and $\mathscr{D}_{2}$.
The proof of this proposition relies on the following lemma, which characterizes the pullback of 11-derivations.
@@
Let $\pi:E\to M$ be a vector bundle equipped with a linear $(1,1)$-tensor field $R\in\Omega^{1}(E,TE)$ covering $r\in\Omega^{1}(M,TM)$, and let $\mathscr{D}=(\nabla,\ell,r)$ be its associated 11-derivation. Let $\mu:P\to M$ be a smooth map and suppose $J\in\Omega^{1}(P,TP)$ is a $(1,1)$-tensor field $\mu$-related to $r$. Then the 11-derivation associated with the pullback $\mu^{*}R$ is given by the triple $\mu^{*}\mathscr{D}=\left(\mu^{*}\nabla,\mu^{*}\ell,J\right)$, whose components are uniquely determined by their action on the pullback of sections:
(37) $(\mu^{*}\nabla)_{v}(\mu^{*}\xi)=\mu^{*}\left(\nabla_{\mathrm{d}\mu(v)}\xi\right),$
(38) $(\mu^{*}\ell)(\mu^{*}\xi)=\mu^{*}(\ell(\xi)),$
@@
Now, applying the correspondence (20) between linear $(1,1)$-tensors and 11-derivations, we directly identify the base component as $J\in\Omega^{1}(P,TP)$. To check the other components, note that the vertical lift of a pullback section $\mu^{*}\xi\in\Gamma(\mu^{*}E)$ at $(p,e)\in\mu^{*}E$ is given by:
@@
On the other hand, since the projection of $(v,w)$ to TPTP is $\mathrm{d}(\mu^{*}\pi)(v,w)=v$, the definition of the connection via the correspondence (20) yields $\left((\mu^{*}\nabla)_{v}(\mu^{*}\xi)\right)^{\uparrow}=\mathscr{L}_{(\mu^{*}\xi)^{\uparrow}}(\mu^{*}R)(v,w)$. Thus, we conclude:
@@
We now proceed to the proof of Proposition 3.17. To this end, we recall the characterization of morphisms of related linear tensor fields established in [undefh, Theorem 2.1]. Let $R_{i}\in\Omega^{1}(E_{i},TE_{i})$ be linear $(1,1)$-tensor fields on vector bundles $E_{i}\to M$ for $i=1,2$, and let $\Phi:E_{1}\to E_{2}$ be a vector bundle map covering the identity. Then
(39) $\mathrm{d}\Phi\circ R_{1}=R_{2}\circ\mathrm{d}\Phi\qquad\iff\qquad\Phi\circ\nabla_{v}^{1}=\nabla_{v}^{2}\circ\Phi,\quad\Phi\circ\ell_{1}=\ell_{2}\circ\Phi\quad\text{and}\quad r_{1}=r_{2},$